OpenStudy (anonymous):

what fraction falls between 3/10 and 3/5 A:5/10 B:8/10 C:4/5 D:2/10

OpenStudy (anonymous):

3/5=6/10 4/5=8/10 Got ur answer?

OpenStudy (anonymous):

Is it "A" "B" "C" or "D"

OpenStudy (anonymous):